OpenStudy (anonymous):

If the density of lead is 11.4 g/ml and the volume is 4.00 ml, what is the mass of a piece of lead?

OpenStudy (mayankdevnani):

\[4.56\times 10^{4}g\]

OpenStudy (mayankdevnani):

@Shiner

OpenStudy (jfraser):

the density of an object is the ratio of its mass to a given volume. Having a density of 11.4g/mL means that 1mL of lead will weigh 11.4g. If you have more than 1mL of lead, it will weigh more than 11.4g. WIth algebra, density calculations look like this:\[D = \frac {mass}{volume}\]In this problem, the density and volume are known, so we can rearrange it to solve for mass instead:\[mass = D*volume\] plugging in values gets:\[mass = 11.4\frac{g}{mL} * 4mL\]See how mL on bottom is cancelled by mL on top? That guarantees that the answer will have the right units, even before looking at the numbers. The answer is 45.6g.

OpenStudy (anonymous):

Thanks guys :)